### v1

# New Perspectives in Online Contract Design: Heterogeneous, Homogeneous, Non-myopic Agents and Team Production

Abstract
This work studies the repeated principal-agent problem from an online learning perspective. The principal’s goal is to learn the optimal contract that maximizes her utility through repeated interactions, without prior knowledge of the agent’s type (i.e., the agent’s cost and production functions).
I study three different settings when the principal contracts with a single agent each round: 1. The agents are heterogeneous; 2. the agents are homogenous; 3. the principal interacts with the same agent and the agent is non-myopic. I present different approaches and techniques for designing learning algorithms in each setting. For heterogeneous agent types, I identify a condition that allows the problem to be reduced to Lipschitz bandits directly. For identical agents, I give a polynomial sample complexity scheme to learn the optimal contract based on inverse game theory. For strategic non-myopic agents, I design a low strategic-regret mechanism. Also, I identify a connection between linear contracts and posted-price auctions, showing the two can be reduced to one another, and give a regret lower bound on learning the optimal linear contract based on this observation.
I also study a team production model. I identify a condition under which the principal’s learning problem can be reformulated as solving a family of convex programs, thereby showing the optimal contract can be found efficiently.
Keywords Online Learning  $\cdot$ Contract Theory
1 Introduction
Contract theory has been a topic of economic research for decades, dating back at least to the seminal work by [Holmström, 1979]. In the contract design problem, a principal (she) interacts with an agent (he) and would like the agent to perform some task on her behalf. However, the exact effort the agent puts into completing the task is unobservable. Therefore, the principal cannot contract upon the action of the agent but rather offer performance-based pay to the agent as an incentive to exert effort. Indeed, this form of principal-agent problem occurs widely in practice. For example, employees receive performance-based bonuses at the end of the year; homeowners selling their houses pay an amount of commission to their agent proportional to the final transaction value.
Most economic literature focuses on exact characterizations of optimal contracts (for a textbook treatment see e.g. [Bolton and Dewatripont, 2004]), thereby assuming the principal has full knowledge of the problem parameters, including the production function and the cost function of the agent. The assumption is far from realistic, and the work by [Ho et al., 2014] is the first to study the problem of learning the optimal contract when the principal andis able to repeatedly interact with agents without any prior knowledge of the agent’s cost function or production function.
The repeated interactions between the principal and agents can be cast as an online learning problem (e.g. [Ho et al., 2014, Zhu et al., 2022]). In repeated interactions, the principal can gradually learn the outcome of the posted contracts and adjust accordingly to find the best contract. This work designs learning algorithms for the principal under different distributional and behavioral assumptions on the agents. Section 1.1 summarizes the results when the principal contracts against a single agent each round. Section 1.2 summarizes the results when the principal contracts against a team of agents each round.
Before giving an overview of the results, I make a brief remark on the modeling choice in this work. I will work with the one-dimensional effort model, i.e., the action space of the agent is a continuous interval on the real line, interpreted as the effort the agent puts into the task. This is in contrast to the discrete action space model which many recent works on contract design from the computer science community have chosen to adopt. While the discrete action space model certainly seems the natural choice when studying contract design from a computational complexity perspective (e.g. [Dütting et al., 2019, Dütting et al., 2021, Dütting et al., 2022, Dütting et al., 2023]), it is puzzling to the author as to why recent papers studying contract design from a machine learning perspective (e.g., [Ho et al., 2014, Zhu et al., 2022]) have also chosen to adopt this discrete action space model. As I will demonstrate, when studying contract design from a machine learning perspective, the one-dimensional effort model is possibly the more natural choice. This modeling choice opens up possibilities for new approaches and techniques to be applied, and allows for more natural assumptions to be made that make the problem more tractable.
1.1 Learning the Optimal Contract: single agent each round
I first study repeated principal-agent problems, where the principal contracts with a single agent in each time period. I study three settings: 1. the agent in each round has different types; 2. the agent in each round has the same type; 3. the principal interacts with the same agent and the agent is non-myopic. Further, I prove a lower bound on the regret for learning the optimal linear contract.
Contracting with hetergeneous agents: continuity conditions and reduction to Lipschitz bandits (Section 3)
The work by [Ho et al., 2014] and [Zhu et al., 2022] studied no-regret algorithms when the principal is contracting against stochastic agent types. They treat the problem as a continuum-armed bandit problem, where the posted contract is the ‘arm’ and the principal can observe the realized reward of the chosen ‘arm’ each round. [Ho et al., 2014] modified the zooming algorithm for Lipschitz bandits, and [Zhu et al., 2022] proposed a discretization algorithm based on spherical codes.
Despite the similarity between online contract design and continuum armed Lipschitz bandits, the previous two works treat the online contract design problem differently and proposed relatively specialized algorithms. However, I show that under a relatively mild condition, the principal’s expected utility, as a function of the contract, becomes (Lipschitz) continuous. Subsequently, the online contract design problem can be directly reduced to Lipschitz bandits.
Contracting against identical agents: a polynomial sample complexity scheme (Section 4)
Next, I study the sample complexity of learning an approximately optimal contract when agents have identical types. Even when faced with identical types, existing works (e.g. [Ho et al., 2014, Zhu et al., 2022, Cohen et al., 2022]) can only achieve exponential sample complexity. This seems to be caused by the nature of their approach, which all model the problem as a continuum-armed bandit problem, in which essentially the principal needs to search for the optimal contract in a high-dimensional space. By contrast, I take an approach based on inverse game theory. At a high level, the agent’s production and cost function can efficiently be recovered by only querying contracts in a low dimensional space (specifically linear contracts). This approach achieves polynomial sample complexity, an exponential improvement over prior works.
The approach is somewhat related to the literature on learning optimal strategies in Stackelberg games. Indeed, the contract design problem can be viewed as a Stackelberg game, where the principal is the leader and the agent is the follower. The principal commits to a contract to which the agent best responds. [Letchford et al., 2009] and [Peng et al., 2019] studied the sample complexity for the leader to learn the optimal strategy to commit to. It is shown in the worst case, the leader requires exponential sample complexity to learn the optimal strategy. Therefore, naively applying their methods would still give an exponential sample complexity in the contract design problem, no better than the existing works (e.g., [Cohen et al., 2022]). However, I show that in the contract design problem, a very natural condition that captures normal economic intuitions (namely the diminishing returns assumption) enables a polynomial sample complexity scheme.
Contracting against a strategic non-myopic agent: a low strategic-regret mechanism (Section 5)
In the above two sets of results, it is assumed the agent is myopic and chooses the action that maximizes his immediate utility each round. However, when the principal repeatedly contracts with the same agent for whom she initially possesses little to no information, the agent can exhibit strategic behaviors and manipulate the principal. For example, the agent may intentionally underperform in low-paying contracts and manipulate the principal into raising the incentive.
This work studies the problem of designing learning mechanisms for contracting against a strategic non-myopic agent, which to the best of my knowledge, is the first work to do so. Note that repeated games with a moral hazard component and uncertainty in the agent’s type have occurred in the economics literature, e.g. the seminal work [Holmström, 1999]. The focus of this line of work in economics is mostly on characterizing the equilibrium behavior. By contrast, our work focuses on designing algorithmic mechanisms for the principal, more similar in spirit to the line of work on pricing against strategic agents (e.g. [Amin et al., 2013, Mohri and Medina, 2015]).
[…]
Learning the optimal linear contract: connection to posted-price auctions and a lower bound (Section 6)
I make an explicit connection between linear contracts (call it contracting) and posted-price auctions (call it pricing). For the contracting problem, the principal aims to find the best linear contract against an agent whose cost is unknown and whose action cannot be perfectly observed. For the pricing problem ([Kleinberg and Leighton, 2003]), the seller aims to find the optimal fixed price for an item against a population of buyers with some unknown demand curve. At first glance, the contracting problem and pricing problem are entirely different. The contracting problem involves a moral hazard problem of hidden action, i.e., the action of the agent cannot be observed. In the pricing problem, the agent’s action can be observed (either a purchased or unpurchased item), but the agent’s private valuation for the item is not observed. I show that in fact, the two problems share a similar structure and can be reduced to one another. Utilizing this connection, I show a lower bound on the regret of the principal for learning the optimal linear contract, via a direct reduction from a result in [Kleinberg and Leighton, 2003].
1.2 Characterizing and Learning the Optimal Contract in Team Production
I also study a team production setting in Section 7, which is a natural extension of the single-agent setting. The seminal work by [Holmstrom, 1982] studied the issue of free-riding and competition in a team production model. More recent works from the computer science community studied a team production model of combinatorial nature from a computational complexity perspective ([Dütting et al., 2023]). By contrast, I study learning algorithms for the principal, which to the best of my knowledge, is the first work to do so.
There is a task that must be completed by a group of $n$ agents. Each agent $i$ can supply an effort $a_{i}\in[0,\infty)$, and the final production is $f(a_{1},\dots,a_{n})$ (assume there is no noise in the production). Further, each agent $i$ has a cost $c(a_{i})$ for exerting effort $a_{i}$. The principal can write contract profiles which are linear, specified by an $n$-tuple $\beta=(\beta_{1},\dots,\beta_{n})$. Under the contract profile $\beta$, agent $i$ receives a share of $\beta_{i}$ of the total outcome, i.e., assuming risk-neutral agents, then agent $i$ has utility
$\beta_{i}f(a_{1},\dots,a_{n})-c(a_{i}).$
The team of agents responds with a Nash equilibrium. Assume $a(\beta)$ is the team’s response to incentive $\beta$, then the principal’s utility, as a function of $\beta$, is
$(1-\sum_{i\in[n]}\beta_{i})f(a(\beta)).$
However, the principal does not have knowledge of the production function $f$, and therefore cannot predict the equilibrium behavior of the team. Further, since the actions of each agent cannot be observed, the only signal the principal can observe (after posting the contract) is the production outcome, i.e., the scalar $f(a(\beta))$. The principal must learn the optimal contract profile with only these limited information.
Consider the following production example with two agents. The production function $f(a,b)=3{a^{1/3}b^{1/3}}$, i.e., a Cobb-Douglas production with decreasing return to scale. The cost function of the two agents are $c_{1}(a)=c_{2}(a)=a$. When the posted contract is $\beta=(\beta_{1},\beta_{2})$, excluding the degenerate equilibrium $(0,0)$, the Nash equilibrium is
$a_{1}=\beta_{1}^{2}\beta_{2},a_{2}=\beta_{2}^{2}\beta_{1}.$
Then the production, expressed as a function of $\beta$, is $f(a(\beta))=3\beta_{1}\beta_{2}$. A first instinct might be to apply some optimization algorithm and solve the following program
$\max_{\beta}(1-\beta_{1}-\beta_{2})(3\beta_{1}\beta_{2}),$
where the objective is the principal’s utility under contract $\beta$. However, it can checked the above objective is not concave, therefore at least the principal can not naively apply some convex optimization algorithm and hope the global optimum can be found. An important observation is that though the utility is not concave, the production $f(a(\beta))=3\beta_{1}\beta_{2}$ is quasiconcave with respect to $\beta$. How does this help us? Instead of optimizing the utility directly, consider the following program
$\min\sum_{i\in[n]}\beta_{i},\quad\text{s.t. }f(a(\beta))\geq k.$
The meaning of the above program is clear: it is the minimum total share that must be distributed to the agents, subject to the constraint that the production must meet a certain threshold $k$. Since $f(a(\beta))$ is quasiconcave, the constraint forms a convex set, and therefore the above program is a convex program! Solving the above program for all production level $k$ gives the optimal contract.
Note however that the quasiconcavity of $f(a(\beta))$ is in fact highly non-trivial and should not be taken for granted. I identify a technical condition that guarantees the quasiconcavity of $f$ with respect to $\beta$. A second note is that the convex program cannot be solved by naive constrained optimization algorithms, such as projected gradient descent, since the constraint set is unknown to the principal. However, recent developments in convex optimization from membership oracles (e.g., [Lee et al., 2018]) exactly suit this setting. The quasiconcavity of $f(a(\beta))$, together with the reformulation of the principal’s problem by the family of convex programs, then gives an efficient learning algorithm for the principal. In particular, it is shown the principal can find an $\varepsilon$-optimal contract using $\widetilde{O}(n^{2}/\varepsilon)$ queries.
[…]
Things to cover:
  • Polynomial sample complexity against a myopic agent
  • The SDFC condition makes the problem instance nice
  • Applying the SDFC condition to stochastic or adversarial agents
  • Lower bound for Linear contracts.
  • Linear contracts with strategic agents
  • General contracts against strategic agents?
  • Reverse Moral Hazard?
2 The Basic Model and Preliminaries
This section describes the basic model when the principal contracts with a single agent, the setting studied in Section 3 to 6. The principal and agent are assumed to be risk-neutral. The one-dimensional effort model consists of the following parameters.
  • $\mathcal{A}=[0,E]$ is the set of actions available to the agent, interpreted as the effort the agent puts into the underlying task. Higher values correspond to higher efforts.
  • $c\mathrel{\mathop{:}}\mathcal{A}\rightarrow\mathbb{R}$ is the cost function of the agent, i.e. $c(a)$ is the cost for exerting effort $a$. $c$ is increasing and convex with $c(0)=0$.
  • $\mathcal{Y}=\mathinner{\left\{\pi_{1},\dots,\pi_{m}\right\}}$ is the production outcome space. There are $m$ possible outcomes. $\pi_{j}$ is the $j$-th level of outcome and assume $0=\pi_{1}<\pi_{2}<\dots<\pi_{m}\leq 1$.
  • $f\mathrel{\mathop{:}}\mathcal{A}\rightarrow\Delta(\mathcal{Y})$ is the production technology of the agent. $f_{j}(a)$ is the probability of obtaining outcome level $j$ when the agent takes action $a$.
  • ${s}\mathrel{\mathop{:}}\mathcal{Y}\rightarrow\mathbb{R}^{+}$: the payment schedule, i.e., the contract. Note that the agent is protected by limited liability in that the payment is non-negative.
It will be sometimes more convenient to work with the complementary distribution function, which is the probability of obtaining an outcome no less than level $j$ when the agent takes action $a$.
$G_{j}(a)\mathrel{\mathop{:}}=\sum_{k=j}^{m}f_{j}(a).$
I will write $u_{\mathsf{A}}(a;{s})$ and $u_{\mathsf{P}}(a;{s})$ to represent respectively the expected utility of the agent and principal when the contract is specified by ${s}$ and the agent takes action $a$. Both parties are risk-neutral. The utility of the agent and principal in this model can then be expressed as:
$u_{\mathsf{A}}(a;s)=\sum_{j\in[m]}s_{j}f_{j}(a)-c(a),$
$u_{\mathsf{P}}(a;s)=\sum_{j\in[m]}(\pi_{j}-s_{j})f_{j}(a).$
For simplicity, I assume the agent’s outside option is 0 and $c(0)=0$, so that it will not hurt for the agent to participate as long as he is protected by limited liability. In a single round of interaction, the principal posts a contract $s$, and the agent takes an action $a$ satisfying the incentive compatibility constraint $u_{\mathsf{A}}(a;s)\geq\max_{a^{\prime}\in\mathcal{A}}u_{\mathsf{A}}(a^{\prime};s)$, i.e., action $a$ gives the agent the highest expected utility among all actions. The goal of the principal is to design the contract $s$ to maximize her utility, captured by the below program
$\max_{s,a}u_{\mathsf{P}}(a;s)$
s.t. $u_{\mathsf{A}}(a;s)\geq u_{\mathsf{A}}(a^{\prime};s)\quad\forall a^{\prime}\in\mathcal{A}.$
Slightly overloading notation, let $u_{\mathsf{P}}(s)\mathrel{\mathop{:}}=\max_{a}u_{\mathsf{P}}(a;s)$, where the maximum is taken over actions $a$ which are incentive compatible under $s$.
In most cases, I will restrict attention to monotone and bounded contracts, as defined below.
Definition 1 (Monotone Contracts).
$s$ is monotone if $s_{j+1}\geq s_{j}$ for any $11\leq j\leq m-$ 1.
Restricting our attention to monotone contracts is very natural, as it prevents the agent from withholding or sabotaging the outcome. This observation has already been made in the important paper by [Innes, 1990].
Definition 2 (Bounded Contracts).
$s$ is bounded if $s_{j}\leq 1$ for any $j\in[m]$.
Though the optimal contract need not be bounded (as in the definition above), imposing an upper bound on the possible payments is natural from a practical perspective. For a more detailed treatment on characterizing the optimal contract with bounded payments, see [Jewitt et al., 2008].
In the following, whenever the term ‘optimal contract’ appears, it should be understood as the optimal contract that is monotone and bounded, unless explicitly stated otherwise.
For any vector $s\in\mathbb{R}^{m}$, denote $\Delta s_{j}=s_{j}-s_{j-1}$, where $s_{0}=0$ is conveniently defined. The following identity is due to Abel’s lemma, which will be used in several places throughout the analysis.
$\sum_{j\in[m]}s_{j}f_{j}(a)=\sum_{j\in[m]}\Delta s_{j}G_{j}(a).$
[…]
2.1 Assumptions
We state our key assumptions on the problem instance below.
Assumption 1.
$c$ is twice differentiable and strongly convex with parameter $\lambda_{0}$.
The second assumption is that the production technology is Lipschitz continuous with respect to the effort level $a$. This is very natural, it implies the production function $f$ cannot change drastically when the agent shifts his effort by a small amount.
Assumption 2.
For any $y\in\mathcal{Y}$, $G_{j}(a)$ is $L$-Lipschitz in $a$ for any $j$.
The third assumption is that the production outcome satisfies diminishing returns with respect to the effort level $a$. Let $r(a)$ be the expected production output of the agent:
$r(a)=\sum_{j\in[m]}\pi_{j}f_{j}(a).$
Assumption 3.
r(a) is increasing, concave, and twice differentiable in $a$.
[…]
remarks TODO
3 Heterogeneous Agents: Continuity Conditions and Reduction to Lipschitz Bandits
I first begin with a discussion on the hardness of computing and characterizing the optimal contract.
Hardness of agent’s optimization problem.
Recall that solving for the optimal contract requires the incentive compatibility constraint:
$u_{\mathsf{A}}(a;s)\geq u_{\mathsf{A}}(a^{\prime};s),\forall a^{\prime}\in\mathcal{A}.$
Notice this is actually a continuum of constraints, and it is unclear how to solve the principal’s program given this set of constraints. It is tempting to replace the above constraint by its first-order condition, so that the principal’s problem becomes
$\max_{s,a}u_{\mathsf{P}}(s;a),\quad\text{s.t. }u^{\prime}_{\mathsf{A}}(a;s)=0,$
where the derivative is taken with respect to $a$. This approach indeed remained popular until the important paper by [Mirrlees, 1999] (circulated earlier in 1975) pointed out it is not always valid. Since then, numerous of economics papers have studied the first-order approach, which seeks to find conditions under which the incentive compatibility constraint can be safely replaced by its first-order condition. Examples include the important papers by [Rogerson, 1985] and [Jewitt, 1988], as well as more recent works [Conlon, 2009, Jung and Kim, 2015]. At a high level, these works essentially finds conditions under which the agent’s optimization problem becomes concave, so that the first-order relaxation is valid.
Hardness of principal’s optimization problem.
In general, the utility of the principal’s utility $u_{\mathsf{P}}(s)$ as a function of $s$ is discontinuous. From a learning perspective, the work by [Ho et al., 2014, Zhu et al., 2022] treats the problem of learning the optimal contract as a continuum-arm bandit problem, however, the discontinuity disables them from using tools from well-understood Lipschitz bandit algorithms, and they proposed modified and relatively specialized discretization-based algorithms for this continuum-arm bandit problem.
3.1 Smoothness of Distribution Function Condition (SDFC)
I introduce a new condition, which I term the smoothness of distribution function condition (SDFC).
Definition 3.
The problem instance is said to satisfy the smoothness of distribution function condition (SDFC) condition if the second derivative of the complementary distribution function is bounded above by the second derivative of the cost function, specifically:
$G^{\prime\prime}_{j}(a)<c^{\prime\prime}(a),\forall j\in[m],a\in\mathcal{A}.$
Remark.
Recall Rogerson’s CDFC requires the function $G$ be concave. Hence, the above SDFC can be seen as a strict relaxation of Rogerson’s CDFC.
Proposition 1.
Under SDFC, for any monotone contract $s$, $u_{\mathsf{A}}(a;s)$ is strictly concave with respect to $a$.
Assumption 4.
The following assumptions are imposed in this section.
  1. 1.
    The SDFC is satisfied with the following:
    $G^{\prime\prime}_{j}(a)\leq\mu<\lambda\leq c^{\prime\prime}(a),\quad\forall j\in[m],a\in\mathcal{A}$
    where the derivative is taken with respect to $a$.
  2. 2.
    $G_{j}(a)$ is $L$-Lipschitz with respect to $a$:
    |$\absolutevalue{G^{\prime}_{j}(a)}\leq L.$
[…]
Remark.
The SDFC is a relaxed and more flexible version than Rogerson’s CDFC. In particular, the first part in the above assumption can be ensured to be satisfied if the principal were able to add a ‘regularizer’ (say of the form $\frac{\lambda_{0}}{2}a^{2}$ for some suitable $\lambda_{0}$) to the agent’s cost function. Practically speaking, this means the principal is artificially increasing the cost function of the agent in favor of more control of the agent’s action. Anecdotal evidence suggests this phenomenon does in fact occur in practice. The idea of regularization in machine learning is well known (see e.g. Chapter 13 in [shalev2014understanding]), but this application to the principal-agent problem seems to be new, and we were not able to find a direct reference. This point possibly merits further research.
Theorem 1.
Under 4, the utility of the principal $u_{\mathsf{P}}(s)$ is Lipschitz (as a function of the contract $s$), specifically:
|$\absolutevalue{u_{\mathsf{P}}(s^{1})-u_{\mathsf{P}}(s^{2})}\leq\left(\frac{4L^{2}}{\lambda-\mu}+2\right).\norm{s^{1}-s^{2}}_{1}.$
3.2 Contracting against hetergeneous agents and reduction to Lipschitz bandits
Let us consider the problem of contracting against heterogeneous agents in a sequential fashion. There is a total of $T$ rounds and in each round, some agent is drawn by nature with whom the principal engages in a contractual relationship. Each agent will be considered as having a private type $\theta$, where the type of the agent represents his production technology and cost function. I consider two scenarios, stochastic agent arrivals and adversarial agent arrivals. In the following, let $u_{\mathsf{P}}(s|\theta)$ denote the utility of contract $s$, when the agent has type $\theta$. Assume that 4 is satisfied for every possible agent type.
Stochastic agent arrival
In the stochastic setting, the agent’s type in each round is drawn from some fixed distribution $\mathcal{D}$ that is unknown to the principal. Then, fixing the type $\theta$, the principal’s utility $u_{\mathsf{P}}(s|\theta)$ is Lipschitz. Consequently $\mathbb{E}_{\theta\sim\mathcal{D}}[u_{\mathsf{P}}(s|\theta)]$ is also Lipschitz. The performance of the principal is measured by the regret, defined as:
$T\cdot\max_{s}\mathbb{E}_{\theta\sim\mathcal{D}}[u_{\mathsf{P}}(s|\theta)]-\sum_{t\in[T]}\mathbb{E}_{\theta_{t}\sim\mathcal{D}}[u_{\mathsf{P}}(s_{t}|\theta_{t})].$
The principal can then directly use the machinery for stochastic Lipschitz bandits (e.g. [Kleinberg et al., 2013]) to optimize her regret.
Proposition 2.
In the stochastic arrival scenario, the principal can achieve a regret $\widetilde{O}(T^{(z+1)/(z+2)})$. Here $z$ is the zooming dimension of the problem instance (which depends on $\mathbb{E}_{\theta\sim\mathcal{D}}[u_{\mathsf{P}}(s|\theta)]$).
Adversarial agent arrival
In the adversarial arrival setting, the agent’s type in each round is chosen by some adversary. The regret of the principal is then defined as:
$\max_{s}\sum_{t\in[T]}u_{\mathsf{P}}(s|\theta_{t})-\sum_{t\in[T]}u_{\mathsf{P}}(s_{t}|\theta_{t}).$
The principal can directly use the machinery for adversarial Lipschitz bandits (e.g. [Podimata and Slivkins, 2020]) to optimize her regret.
Proposition 3.
In the adversarial arrival scenario, the principal can achieve a regret $\widetilde{O}(T^{(z+1)/(z+2)})$. Here $z$ is the adversarial zooming dimension of the problem instance (which depends on the sequence of functions $u_{\mathsf{P}}(s|\theta_{1}),\dots,u_{\mathsf{P}}(s|\theta_{T})$).
Remark.
It seems important to point out the conceptual differences between this approach and approaches in prior work (e.g. [Ho et al., 2014, Zhu et al., 2022]). While prior works design specialized discretization-based algorithms for the online contract design problem, the approach in this work finds conditions under which the problem can directly be reduced to Lipschitz bandits. The condition is based on the first-order approach (e.g. [Rogerson, 1985]), and therefore either an advancement in the first-order approach or an advancement in Lipschitz bandits will likely lead to an improvement of the approach presented here. Also, I believe the reduction also works under possibly more relaxed conditions. Specifically, relaxing 4 may lead to a more relaxed continuity condition on the utility function $u_{\mathsf{P}}(s)$ (e.g. Holder continuity instead of Lipschitz continuity), and then a reduction to continuum armed bandits with mitigated Lipschitz conditions (e.g. [Locatelli and Carpentier, 2018]) may be possible.
4 Identical Agents: a Polynomial Sample Complexity Scheme
Algorithm 1 Polynomial Sample Complexity for Identical Agents
[…]
Input: $\varepsilon,\delta$. Algorithm returns $\varepsilon$-approximately optimal contract with probability $1-\delta$.
Set $\varepsilon_{c}=\frac{\lambda\varepsilon}{2L^{2}},N=\frac{2\log(\beta_{\max}/(\delta\varepsilon_{c}))}{\varepsilon^{2}}$
Set $\mathcal{C}_{D}=\mathinner{\left\{0,1,\dots,\left\lfloor\beta_{\max}/\varepsilon_{c}\right\rfloor\right\}}\cdot\varepsilon_{c}$, the discretized contract space
For each contract $\beta^{(i)}\in\mathcal{C}_{D}$, the principal queries this contract $N$ times
Principal records the empirical complementary CDF $\hat{G}_{j}(a^{(i)})$
Principal computes the empirical average outcome using $\hat{G}$:
$\hat{r}(a^{(i)})=\mathbb{E}_{j\sim\hat{G}(a^{(i)})}[\pi_{j}]$
Principal compute the confidence interval for the cost as:
${c}^{\mathtt{LCB}}(a^{(i)})$ $\mathrel{\mathop{:}}=\sum_{j=1}^{i}\beta^{(j-1)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))-3\varepsilon,$
${c}^{\mathtt{UCB}}(a^{(i)})$ $\mathrel{\mathop{:}}=\sum_{j=1}^{i}\beta^{(j)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))+3\varepsilon$
Record the estimated parameters $\hat{r}(a^{(i)}),\hat{G}_{j}(a^{(i)}),{c}^{\mathtt{UCB}}(a^{(i)}),{c}^{\mathtt{LCB}}(a^{(i)})$ for all $i\in\mathinner{\left\{0,1,\dots,\left\lfloor 1/\varepsilon_{c}\right\rfloor\right\}}$
Solve a family of Linear Programs $\mathtt{LP}(a^{(i)})$ and return the approximately optimal contract (specified in Section 4.2)
In this section, I study the sample complexity of learning an approximately optimal contract when the agent in each round has identical types. I.e., each agent has the same cost function and production function. Note this is essentially a special case of the previous section. While one can apply the previous approach and reduce the problem to Lipschitz bandits, I show a different approach based on inverse game theory that achieves $O({poly}(1/\varepsilon))$ sample complexity for learning the $\varepsilon$-optimal contract when the agents have identical types. Here, a $\varepsilon$-optimal contract is defined as a contract achieving utility no less than the optimal contract minus $\varepsilon$.
The problem of finding the optimal contract can be viewed as a Stackelberg game, where the principal is the leader and the agent is the follower who best responds to the principal’s action. The previous works [Letchford et al., 2009, Peng et al., 2019] study the sample complexity needed for the leader to find the optimal strategy to commit to. However, it is shown in the worst case, the leader will need exponentially many samples to learn the optimal strategy. These prior works essentially show that, if for each action of the follower, the leader can find some strategy profile that implements it, then she can efficiently recover the follower’s utility matrix and find the optimal strategy to commit to. However, it is also shown that, there may exist some action of the follower, such that finding some strategy profile that implements it may take exponential sample complexity.
Our goal is to apply the appropriate economic intuitions in the contract design problem and circumvent this exponential sample complexity. It is shown under the very natural diminishing returns assumption, that the agent’s response will be well-behaved under linear contracts. Specifically, the agent’s response will be continuously increasing as the linear contract increases.
Assumption 5.
The following assumptions are made in this section.
  1. 1.
    $c(a)$ is twice-differentiable and $\lambda$-strongly convex.
  2. 2.
    $G_{j}(a)$ is $L$-Lipschitz.
  3. 3.
    $r(a)$ is strictly increasing, concave, and twice differentiable in $a$.
  4. 4.
    The maximum level of effort $E\leq\frac{L}{\lambda}$ and can be implemented by the linear contract $\beta_{\max}$.
Remark.
The third part of the assumption amounts to the diminishing returns intuition. The fourth part is without loss of generality, since the cost of implementing any effort level greater than $\frac{L}{\lambda}$ already outweighs the expected outcome, hence cannot be optimal to the principal. In studying the sample complexity, I will mostly treat $\lambda,L,\beta_{\max}$ as constants and focus on the dependence on $\varepsilon$.
For each contract linear contract $\beta$, if the principal queries this contract enough times, then she can effectively recover the production and cost at the point $a(\beta)$ in the agent’s action space. By appropriately discretizing the linear contract space $[0,\beta_{\max}]$ and querying each contract in the discretized space enough times, the principal essentially recovers the whole production function and cost function of the agent (up to some small error) and can use a family of linear programs to find the approximately optimal contract. Note that $\beta_{\max}$ may be greater than 1. This means the principal may need to use linear contracts that give the agent more than the actual outcome to implement some actions in order to learn the production function. This is a minor drawback of the proposed algorithm.
4.1 Recovering the agent’s production and cost function with linear contracts
In the first stage of the algorithm, the principal queries linear contracts and learns the relevant parameters. The principal will use the observed outcomes to estimate the production technology. Let $a(\beta)$ be the agent’s best response under $\beta$. The principal will discretize the linear contract space $[0,\beta_{\max}]$ into $\varepsilon_{c}$-bins. Specifically, the principal will query contracts from the set $\mathcal{C}_{D}$, where:
$\mathcal{C}_{D}=\mathinner{\left\{0,1,\dots,\left\lfloor\beta_{\max}/\varepsilon_{c}\right\rfloor\right\}}\cdot\varepsilon_{c}.$
Let $a^{(i)}=a(i\cdot\varepsilon_{c})$ for $0\leq i\leq\left\lfloor\beta_{\max}/\varepsilon_{c}\right\rfloor$. I.e., $a^{(i)}$ is the response of the agent under the contract $\beta=i\cdot\varepsilon_{c}$. Denote
$\mathcal{A}_{D}=\mathinner{\left\{a^{(i)}\mathrel{\mathop{:}}0\leq i\leq\left\lfloor\beta_{\max}/\varepsilon_{c}\right\rfloor\right\}}.$
It should be noted that the exact value of $a^{(i)}$ (and by extension the elements in $\mathcal{A}_{D}$) is unknown. Rather, the principal’s goal is to obtain accurate estimates of the production function $G_{j}(a^{(i)})$ and the cost of the agent $c(a^{(i)})$. The exact identity of $a^{(i)}$ is unimportant.
The principal computes the confidence interval for the cost and probability distribution of the induced action $a$ under $\beta$. The algorithm achieves accurate estimations, as in the next theorem.
Theorem 2.
With probability $1-\delta$, for any $a^{(i)}\in\mathcal{A}_{D},j\in[m]$, the error of the estimated parameters in Algorithm 1 can be bounded as:
$G_{j}(a^{(i)})$ $\in\left[\hat{G}_{j}(a^{(i)})-\varepsilon,\hat{G}_{j}(a^{(i)})+\varepsilon\right]$
$c(a^{(i)})$ $\in[{c}^{\mathtt{LCB}}(a^{(i)}),{c}^{\mathtt{UCB}}(a^{(i)})].$
Moreover, ${c}^{\mathtt{UCB}}(a^{(i)})-{c}^{\mathtt{LCB}}(a^{(i)})\leq 9\varepsilon$.
4.2 Computing the approximately optimal contract
The second stage of the algorithm involves solving a family of linear programs. It is known that the optimal contract can be solved using a family of linear programs ([Grossman and Hart, 1992]), here the principal is essentially using the same family but with the estimated parameters (instead of the true parameters, which are unknown) obtained in the first stage.
Definition 4.
Define
$u^{\mathtt{UCB}}_{\mathsf{A}}(a^{(i)};s)\mathrel{\mathop{:}}=\sum_{j\in[m]}\Delta s_{j}G^{\mathtt{UCB}}_{j}(a^{(i)})-c^{\mathtt{LCB}}(a^{(i)})$
$u^{\mathtt{LCB}}_{\mathsf{A}}(a^{(i)};s)=\sum_{j\in[m]}\Delta s_{j}G^{\mathtt{LCB}}_{j}(a^{(i)})-c^{\mathtt{UCB}}(a^{(i)}).$
For each $a^{(i)}\in\mathcal{A}_{D}$, the principal solves the following linear program, denote the program by $\mathtt{LP}(a^{(i)})$ and its value by $\mathtt{APX}(a^{(i)})$.
$\max_{s}\quad$ $\hat{r}(a^{(i)})-\sum_{j\in[m]}s_{j}\hat{f}_{j}(a^{(i)})$
$u_{\mathsf{A}}^{\mathtt{UCB}}(a^{(i)};s)$ $\geq u_{\mathsf{A}}^{\mathtt{LCB}}(a^{(k)};s)-\varepsilon\quad\forall a^{(k)}\in\mathcal{A}_{D}$
$s_{j}$ $\geq s_{j-1}$
$s_{j}$ $\geq 0$
Definition 5.
Given contract $s$, action $a$ is $\delta$-IC under $s$ if for any $a^{\prime}\in\mathcal{A}$:
$u_{\mathsf{A}}(a;s)\geq u_{\mathsf{A}}(a^{\prime};s)-\delta.$
A contract-action pair $(s,a)$ is said to be $\delta$-IC if $a$ is an $\delta$-IC under $s$.
Using the family of linear programs with estimated parameters, the principal can find a $\delta$-IC contract that approximates the utility of the optimal contract.
Proposition 4.
The family of linear programs give a $22\varepsilon$-IC contract that is $6\varepsilon$-optimal.
Then, using the very interesting observation in [Dütting et al., 2021], which allows the principal to convert any $\delta$-IC contract into an exact-IC contract with a small loss in utility, we arrive at the following theorem.
Theorem 3.
The principal can find a $6\varepsilon+10\sqrt{\varepsilon}$-optimal contract using $\widetilde{O}(1/\varepsilon^{3})$ samples.
Corollary 1.
The principal can find a $\varepsilon$-optimal contract using $\widetilde{O}(1/\varepsilon^{6})$ samples.
Remark.
Note that the sample complexity is independent of the number of outcomes $m$. As a consequence, our scheme can be adapted to the case where the outcome space is continuous, e.g. $\mathcal{Y}=[0,Y]$, a continuous interval. Specifically, this involves discretizing the outcome space and applying our algorithm on the discretized outcome space. Bounding the discretization error in the outcome space requires some additional regularity conditions on the production function. The exact details are omitted.
5 Non-myopic agents: a low strategic-regret mechanism
Algorithm 2 Contracting against non-myopic agents using delayed elimination
Input: $\mathcal{C}$ (contract space), $T$ (time horizon), $D$ (delay parameter), $\delta$ (perturbation parameter)
Principal maintains active set of arms $A$, initialized as $\mathcal{C}$
while $t<T$ do
     Pull each arm $\beta\in A$ and observe outcome $y_{t}$
     For each arm $\beta\in A$, update the adjusted UCB and LCB index:
$S$ =$=\mathinner{\left\{1\leq\tau<t-D\mathrel{\mathop{:}}\beta_{\tau}=\beta\right\}}$
$\hat{u}(\beta)$ $=\sum_{\tau\in S}y_{\tau}/\absolutevalue{S}$ |
$\mathtt{UCB}(\beta)$ $=\hat{u}(\beta)+\sqrt{2\log T/\absolutevalue{S}}+\delta$
$\mathtt{LCB}(\beta)$ $=\hat{u}(\beta)-\sqrt{2\log T/\absolutevalue{S}}-\delta,$
     $|t\mathrel{\mathop{:}}=t+\absolutevalue{A}$ |
     For each $\beta\in A$, if $\mathtt{UCB}(\beta)<\max_{\beta^{\prime}\in A}\mathtt{LCB}(\beta^{\prime})$, remove $\beta$ from $A$
end while
This section studies the problem of contracting against a strategic agent. I first begin with the problem description. There is a single principal and a single agent. The principal repeatedly interacts with the agent. The agent is strategic, in the sense that he is optimizing his long-term utility instead of acting myopically and best-responding each round. The goal of the principal is to design a mechanism to maximize her own utility against a non-myopic agent.
I will mostly focus on linear contracts, in other words, the principal will be able to post commission-type contracts parameterized by $\beta\in[0,1]$. Notice that when focusing on linear contracts, for the risk-neutral principal and agent, only the expected outcome for each action $a$ matters, rather than the exact specification of the production technology $f$. Hence, let us simplify our model further and assume that when the agent supplies effort $a$, the expected outcome is a subgaussian random variable centered at $a$. Note that this is without loss of generality since one can always perform an appropriate reparameterization on the agent’s action space. For simplicity, assume that the output is bounded in $[0,1]$ by performing an appropriate normalization.
Now, the interaction can be specified as follows.
  1. 1.
    At the beginning of the interaction, the principal announces her mechanism $\mathcal{M}$. Then, the agent decides on his strategy $\mathcal{S}$, which achieves highest expected utility under the mechanism $\mathcal{M}$.
  2. 2.
    At each round $t\in[T]$, the principal posts the linear contract specified by $\beta_{t}$ according to her mechanism $\mathcal{M}$.
  3. 3.
    The agent supplies an effort level $a_{t}\in[0,E]$, incurs cost $c(a_{t})$.
  4. 4.
    Both parties observe the production output $y_{t}$, a subgaussian random variable centered at $a_{t}$. The interaction proceeds to the next round (go to step 2).
The principal’s mechanism is a family of functions $\mathcal{M}_{1\mathrel{\mathop{:}}T}$:
$\mathcal{M}_{t}\mathrel{\mathop{:}}(\beta,y)_{1\mathrel{\mathop{:}}t-1}\rightarrow\Delta(\mathcal{C}).$
In other words, the principal decides the contract at round $t$ based on all observed information so far, this includes the posted contracts as well as the realized outcomes.
The principal announces her mechanism at the start of the interactions. Given the mechanism of the principal, the agent chooses a strategy maximizing his discounted expected total utility. The agent’s strategy is a family of functions:
$\mathcal{S}_{t}\mathrel{\mathop{:}}(\beta_{t},(\beta,y,a)_{1\mathrel{\mathop{:}}t-1},\mathcal{M})\rightarrow\Delta(\mathcal{A}).$
In other words, the agent observes the current contract $\beta_{t}$ and the history of $(\beta,y,a)_{1\mathrel{\mathop{:}}t-1}$, and decides on an effort level. Define $u_{\mathsf{A},t}\mathrel{\mathop{:}}=\beta_{t}y_{t}-a_{t}$ as the agent’s immediate utility in round $t$. Then the agent chooses a strategy to maximize his total expected discounted utility:
$\max_{\mathcal{S}}\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t}u_{\mathsf{A},t}\right].$
where the expectation should be understood as taken with respect to randomness in the production function and possible randomness in the principal’s mechanism and agent’s strategy.
Let $u_{\mathsf{P},t}\mathrel{\mathop{:}}=(1-\beta_{t})y_{t}$ denote the principal’s utility in round $t$. Let $u(\beta)\mathrel{\mathop{:}}=(1-\beta)a(\beta)$ denote the expected utility to the principal under the linear contract $\beta$, assuming the agent responds truthfully.
The performance of the mechanism is measured in terms of the strategic regret, defined as the difference between the utility using the best linear contract and the utility obtained using the mechanism:
$\mathtt{Reg}(\mathcal{M})=\max_{\beta\in\mathcal{C}}Tu(\beta)-\mathbb{E}\left[\sum_{t\in[T]}u_{\mathsf{P},t}\right].$
I first assume there is a finite set of contracts the principal can choose from. Then I extend the analysis to $\mathcal{C}=[0,1]$, and also when $\mathcal{C}$ can be any monotone and bounded contracts.
In the following, use $a(\beta_{t})$ to denote the truthful (i.e. myopic best-response) at round $t$, and use $\delta_{t}\mathrel{\mathop{:}}=a_{t}-a(\beta_{t})$ to denote the deviation from the myopic best-response. Also define $T_{\gamma}$ as $T_{\gamma}\mathrel{\mathop{:}}=\sum_{t=1}^{T}\gamma^{t}.$ In other words, $T_{\gamma}$ is an upper bound on the maximum possible utility that the agent can receive starting from the beginning. This section assumes the following.
Assumption 6.
$c(a)$ is $\lambda$-strongly convex.
5.1 Algorithm for finite set of candidate contracts
Let $\mathcal{C}=\mathinner{\left\{\beta^{(1}),\dots,\beta^{(K)}\right\}}$ be the set of candidate contracts. The algorithm is based on the idea of delayed elimination, adapted from [Haghtalab et al., 2022]. , Each contract is treated as an arm in a multi-arm bandit problem. The principal maintains a set of active arms and computes adjusted UCB and LCB indices using only information from $D$ rounds ago. The indices are further adjusted with a perturbation parameter $\delta$. Arms are eliminated if the UCB index for some arm is lower than any other arm’s LCB index.
Definition 6 (Deviation Cost).
The agent’s deviation cost at round $t$ is defined as:
$u_{\mathsf{A}}(a(\beta_{t});\beta_{t})-u_{\mathsf{A}}(a_{t};\beta_{t}).$
Lemma 1.
The agent’s immediate deviation cost at round $t$ can be lower bounded by $\lambda\delta_{t}^{2}$.
Proof.
Recall $\beta_{t}$ is the posted contract at round $t$ and $\delta_{t}=a(\beta_{t})-a_{t}$. Under the contract $\beta_{t}$, let $u_{\mathsf{A}}(a)=\beta_{t}a-c(a)$ be the expected utility for the agent when taking action $a$. Then $u_{\mathsf{A}}(a)$ is strongly concave with respect to $a$. The lemma then follows by strong concavity:
$u_{\mathsf{A}}(a(\beta_{t}))-u_{\mathsf{A}}(a_{t})\geq\lambda\delta_{t}^{2}.\qed$
Lemma 2.
Suppose the agent is using his optimal strategy. Then at each round $t$, $\lambda\delta_{t}^{2}\leq\gamma^{D}\cdot T_{\gamma}$.
Proof.
The action of the agent at round $t$ can only affect the decisions of the principal after $D$ rounds by the delayed feedback nature of the algorithm. The maximum utility that can be collected after $D$ rounds can be upper bounded by:
$E\cdot\sum_{\tau=t+D}^{T}\gamma^{\tau}\leq\gamma^{t+D}\cdot T_{\gamma}.$
The immediate discounted deviation cost cannot be greater than the maximum possible utility that can be collected after $D$ rounds, hence:
$\lambda\gamma^{t}\delta_{t}^{2}\leq\gamma^{t+D}\cdot T_{\gamma}.$
Dividing both sides by $\gamma^{t}$ completes the proof. ∎
The following lemma bounds the regret of a $D$-delayed elimination algorithm with $\delta$-bounded perturbation. Denote $\Delta_{i}\mathrel{\mathop{:}}=\max_{\beta\in\mathcal{C}}u_{\mathsf{P}}(\beta)-u_{\mathsf{P}}(\beta^{(i)})$.
Lemma 3 ([Haghtalab et al., 2022]).
The regret of $\mathtt{DelayedElimination}$ can be bounded by $O(\sum_{\Delta_{i}>0}\frac{\log T}{\Delta_{i}}+\delta T+D\log K)$.
The below theorem follows by choosing the appropriate parameters.
Theorem 4.
Using Algorithm 2, the strategic regret can be bounded by $O(\sum_{\Delta_{i}>0}\frac{\log T}{\Delta_{i}}+\frac{\log(TT_{\gamma}/\lambda)\cdot\log K}{\log(1/\gamma)})$.
Proof.
Choose the parameter $D=\frac{\ln(T^{2}T_{\gamma}/\lambda)}{\ln(1/\gamma)}$. Then by lemma 2, the deviation each round can be bounded by $1/T$. The theorem then follows immediately by the previous lemma. ∎
5.2 Extending the contract space
All linear contracts
Assume $\mathcal{C}=[0,1]$, i.e., the principal can posts any linear contract. Our general strategy will be to discretize the space and use the algorithm from the previous section. I make an assumption similar to [Kleinberg and Leighton, 2003].
Assumption 7.
Assume there exists a unique $\beta^{*}\in[0,1]$, such that $u(\beta^{*})>u(\beta^{\prime})$ for any $\beta^{\prime}\neq\beta$ and that $u^{\prime\prime}(\beta^{*})>0$.
Theorem 5.
There exists some algorithm that achieves $O(\sqrt{T\log T}+\frac{\log(TT_{\gamma}/\lambda)\log(T/\log T)}{\log(1/\gamma)})$ regret.
Proof.
We shall discretize the contract space $[0,1]$ into bins, specifically let $\mathcal{C}_{K}=\mathinner{\left\{\frac{1}{K},\dots,\frac{K}{K}\right\}}$. Let $\beta^{*}$ be the optimal contract in the discretized space. The discretization error through $T$ rounds can be bounded by $O(T/K^{2})$, and the inverse sum can be bounded by $O(K^{2})$ (see Lemma 4.1 and Lemma 4.2 in [Haghtalab et al., 2022], which are restatements of the original proof in [Kleinberg and Leighton, 2003]). Therefore the total regret is:
$O(K^{2}\log T+\frac{T}{K^{2}}+\frac{\log(TT_{\gamma}/\lambda)\log K}{\log(1/\gamma)}).$
Choosing $K^{2}=\sqrt{T\log T}$ gives the final bound. ∎
Any monotone and bounded contract
Assume $\mathcal{C}$ is the set including any monotone and bounded contract. In this case, the technique in Section 4 (for designing a polynomial sample complexity scheme) can be combined with the idea of using delayed feedback in this section. Recall the algorithm from Section 4 first requires the principal to post linear contracts to learn the agent’s production and cost function. After obtaining the observations, the principal should wait for an appropriate number of rounds before implementing the computed approximately optimal contract. Again, the delay ensures the agent’s response does not deviate too much from his truthful best response, and therefore the estimated production and cost functions are accurate.
6 Connection to Posted-Price Auction and a Lower Bound
The analysis in the previous section seem to bear some similarity to that from the dynamic pricing literature (e.g. [Kleinberg and Leighton, 2003]). In this section, I make the connection more explicit and, utilizing the connection, prove a lower bound on the regret for learning the optimal linear contract.
Consider the contracting problem with a binary outcome setting, i.e. $\mathcal{Y}=\mathinner{\left\{0,1\right\}}$. An outcome of 1 indicates the success of the task and 0 indicates failure. The action space of the agent is $[0,1]$. When the agent takes action $a$, the success probability is $a$ and the cost to the agent $c(a)$. The expected utility under linear contract $\beta$ is then $(1-\beta)a(\beta)$. Let us make the change of variable $\upalpha=1-\beta$, i.e. $\upalpha$ is the fraction of the outcome the principal keeps for herself. Then overloading notation letting the agent’s best response be $a(\upalpha)$ under $\upalpha$, the expected utility under $\upalpha$ is:
$\upalpha\cdot a(\upalpha).$
Note that by convexity of $c(\cdot)$, the function $a$ is monotone decreasing with respect to $\upalpha$. I shall refer to $a(\cdot)$ as the response curve (under cost function $c$).
Now recall the pricing problem as specified in [Kleinberg and Leighton, 2003]. The buyer has a valuation drawn from the demand curve $D(\cdot)$, where $D(p)$ is the probability that the buyer makes a purchase when the price is $p$. In general, the demand curve $D(\cdot)$ should be monotone decreasing. The expected utility of the seller on posting price $p$ is then
$p\cdot D(p).$
There is now a direct relation between the contracting problem and the pricing problem: when the demand curve $D$ and the response curve $a$ take the same shape, the two problems in fact become the same. To elaborate, for any $\upalpha=p$, the principal observes a binary signal with mean $a(\upalpha)$ in the contracting problem, which is exactly the same distribution as the dynamic pricing problem; and the expected utility are exactly same in both problems. Further note that in the contracting problem, given a contract $\beta$, the agent maximizes $\beta a-c(a)$, and the first order condition gives that $c^{\prime}(a)=\beta$. Hence the response curve $a(\cdot)$ is essentially the inverse of $(c^{\prime})^{-1}$.
Now consider the problem of learning the optimal linear contract against a (myopic) agent who best responds each round, for a total of $T$ rounds. Using the above observation, one can prove a lower bound on the regret by directly using a result in [Kleinberg and Leighton, 2003]. Here, the regret is defined as the difference between the total utility collected by the principal and the maximum total utility of the best fixed linear contract.
Theorem 6.
Consider the binary outcome setting as described above and suppose the principal must post linear contracts in $[0,1]$. There exists a distribution of cost functions such that when the agent has a cost function randomly sampled from the distribution, the principal incurs expected regret $\Omega(\sqrt{T})$ for any possible contracting strategy she may adopt.
Proof.
[Kleinberg and Leighton, 2003] studied the pricing problem, and they constructed a distribution $\mathcal{D}$ over demand functions, such that any algorithm will suffer expected regret $\Omega(\sqrt{T})$ against a randomly drawn demand function from $\mathcal{D}$. Further, it can be shown in their construction, that each $D$ in the support of $\mathcal{D}$ is strictly decreasing. For each demand curve $D$, construct a cost function $c$ as follows:
$c(a)=\int_{x=0}^{a}[1-D^{-1}(x)]\operatorname{d\!}x.$
Then the response curve $a(\cdot)$ under $c$ will exactly coincide with $D(\cdot)$. Any algorithm for the contracting problem then suffers the same regret as the pricing problem, immediately implying a $\Omega(\sqrt{T})$ lower bound for the contracting problem. ∎
Remark.
Note that a lower bound of order $\Omega(T^{2/3})$ also appeared in [Zhu et al., 2022]. There is no contradiction, as there is a difference in the modeling choice. While they assumed the action space of the agent is discrete, this work assumed the one-dimensional effort model. Further, their lower bound is based on a somewhat contrived construction, while the approach here observes and exploits a connection to posted-price auctions, which allows us to directly use a result in [Kleinberg and Leighton, 2003].
7 Team Production
This section departs from the single-agent setting and studies the problem of characterizing and learning the optimal contract in a team production setting. I start with a description of the model.
The principal contracts with a group of $n$ agents. When each agent $i$ takes an action $a_{i}\in[0,\infty)$, the production is $f(a)$, where $a$ denotes vector $a=(a_{1},\dots,a_{n})$. For simplicity, assume there is no noise in the production. Also assume the cost function of each agent $c(a)=a$. Note that this is without loss of generality, since for any increasing and continuous cost function $c$ with $c(0)=0$, one can perform an appropriate reparameterization so that $c(a)=a$.
The principal can write linear contracts specified by a $n$-tuple $\beta=(\beta_{1},\dots,\beta_{n})$. In this section, I will use $\beta_{i}$ to denote agent $i$’s share under contract profile $\beta$. Assuming the agents are risk-neutral, the utility of agent $i$ under contract profile $\beta$ and the teams action $a$ is then:
$\beta_{i}f(a)-a_{i}.$
It is assumed that each agent knows the production function so that they will respond with the equilibrium action profile $a=(a_{1},\dots,a_{n})$ that satisfies the following:
$a_{i}\in\arg\max_{a^{\prime}_{i}}\beta_{i}f(a_{-i},a^{\prime}_{i})-a^{\prime}_{i}.$
I restrict attention to the equilibrium that satisfies the first-order conditions. To do so, I make the following assumption, which is relatively standard.
Assumption 8.
$f$ is strictly concave, continuously differentiable, and satisfies the Inada conditions.
Lemma 4.
For any $\beta$, there exists a unique equilibrium satisfying the first-order conditions:
$\beta_{i}\partial_{i}f(a)=1.$
Proof.
Without loss of generality assume $\beta>0$. We need to show the equation
$\nabla f(a)=(1/\beta_{1},\dots,1/\beta_{n})$
has a unique solution. Uniqueness is implied by the strict concavity of $f$. Existence is implied the the Poincare-Miranda theorem with the Inada conditions. ∎
When the contract is $\beta$ and the agents respond with the action profile $a$, the principal enjoys utility
$(1-\sum_{i}\beta_{i})f(a).$
The interaction protocol in each time period is then summarized as follows.
  1. 1.
    The principal posts contract $\beta$.
  2. 2.
    The agent responds with the equilibrium action profile $a$ satisfying the first-order conditions.
  3. 3.
    The principal observes the outcome $f(a)$ and distributes the promised shares, herself enjoying utility $(1-\sum_{i}\beta_{i})f(a)$. Further, the principal does not observe the action profile $a$.
The goal of the principal is to find the (approximately) optimal contract profile using a small number of queries (each round counts as a query). In particular, I show the principal can find an $\varepsilon$-optimal contract profile using $\widetilde{O}(n^{2}/\varepsilon)$ queries.
7.1 A sufficient condition for quasiconcavity
The main goal of this subsection is to provide a sufficient condition for which the production $f(a(\beta))$, as a function of $\beta$, is quasiconcave.
Assumption 9.
$f$ is explicitly additive so that there exists functions $h,g_{1\mathrel{\mathop{:}}n}$ such that $f(a)=h(\sum_{i\in[n]}g_{i}(a_{i}))$, where $h$ is a monotonic transformation, i.e.,
$\sum_{i\in[n]}g_{i}(a_{i})>\sum_{i\in[n]}g_{i}(b_{i})\Rightarrow h(\sum_{i\in[m]}g_{i}(a_{i}))>h(\sum_{i\in[m]}g_{i}(b_{i})),$
and $g_{i}$ is strictly increasing and concave for each $i$. Further the function
$y_{i}(\cdot)=g_{i}\circ(1/g^{\prime}_{i})^{-1}(\cdot)$
is well-defined, strictly increasing, and concave.
[…]
Lemma 5.
$g(\beta)=\sum_{i\in[n]}y_{i}(\beta_{i}h^{\prime}(g(\beta)))$.
Proof.
In this proof denote $a=a(\beta)$. Since $a$ forms an equilibrium and the first-order conditions are met:
$\beta_{i}=\frac{1}{h^{\prime}(\sum_{i}g_{i}(a_{i}))}\cdot\frac{1}{g^{\prime}_{i}(a_{i})}.$
Noting
$g_{i}(a_{i})=y_{i}(1/g^{\prime}_{i}(a_{i}))=y_{i}(\beta_{i}h^{\prime}(g(\beta)))$
then for any $\beta$,
$g(\beta)=\sum_{i}y_{i}(\beta_{i}h^{\prime}(g(\beta))).\qed$
Lemma 6.
Fix any $t_{0}$. Denote $S_{1}\mathrel{\mathop{:}}=\mathinner{\left\{\beta\mathrel{\mathop{:}}g(\beta)\geq t_{0}\right\}}$. Then $S_{1}$ is a convex set.
Proof.
Denote $S_{2}\mathrel{\mathop{:}}=\mathinner{\left\{\beta\mathrel{\mathop{:}}t_{0}\leq\sum_{i}y_{i}(\beta_{i}h^{\prime}(t_{0}))\right\}}$. We shall show $S_{1}=S_{2}$.
We first show $S_{2}\subset S_{1}$. Take any element $\beta\in S_{2}$, then we need to show $g(\beta)\geq t_{0}$. For sake of contradiction assume $g(\beta)=t_{1}<t_{0}$. Then
$t_{1}=\sum y_{i}(\beta_{i}h^{\prime}(t_{1})).$
However, there must exists a $\zeta=(\zeta_{1},\dots,\zeta_{n})>\beta$, such that $g(\zeta)=t_{0}$. Then
$t_{0}=\sum y_{i}(\zeta_{i}h^{\prime}(t_{0}))>\sum y_{i}(\beta_{i}h^{\prime}(t_{0})),$
which is a contradiction with the fact that $\beta\in S_{2}$.
We next show $S_{1}\subset S_{2}$. Take any $\beta\in S_{1}$, then we need to show
$t_{0}\leq\sum y_{i}(\beta_{i}h^{\prime}(t_{0})).$
Since $t_{0}\leq t_{1}$, there must exists a contract $\zeta=(\zeta_{1},\dots,\zeta_{n})$ such that $\zeta<\beta$ and that $g(\zeta)=t_{0}$, therefore
$t_{0}=\sum_{i}y_{i}(\zeta_{i}h^{\prime}(t_{0}))\leq\sum_{i}y_{i}(\beta_{i}h^{\prime}(t_{0})).$
Since the set $S_{2}$ is convex, $S_{1}$ is convex. ∎
Theorem 7.
$f(a(\beta))$, as a function of $\beta$, is quasiconcave.
Proof.
Note $f(a(\beta))=h(g(\beta))$. By the previous Lemma, $g$ is quasiconcave. $f$ is then a composition of an increasing function with a quasiconcave function, therefore it is quasiconcave. ∎
Example 1.
9 is satisfied for the CES production family when the substitution parameter is non-positive. Note that non-positive substitution parameter ensures that every agent is ‘essential’, i.e., $f(a)=0$ if $a_{i}=0$ for some $i$.
  1. 1.
    Consider the CES production function with negative substitution parameter. I.e.,
    $f(a)=(\sum_{i}\rho_{i}a_{i}^{-k})^{\nicefrac{{-d}}{{k}}}.$
    where $k>0$. Then we can define $h(x)=-x^{\nicefrac{{-d}}{{k}}}$ and $g_{i}(a_{i})=-\rho_{i}a_{i}^{-k}$. Then $f(a)=h(\sum g_{i}(a_{i}))$.
  2. 2.
    Consider the Cobb-Douglas production function: $f(a)=\prod_{i}a_{i}^{k_{i}}$, which is a special case of the CES production function as the substitution parameter approaches 0. Then $f(a)$ can be expressed in explicitly additive form $f(a)=\exp(\sum_{i}k_{i}\ln a_{i})$.
In both cases, it can be verified that 9 is verified.
7.2 Characterizing and Learning the Optimal Contract
Algorithm 3 Finding the Optimal Contract in Team Production
$k=0$
$\mathtt{OptUtil}=0$, $\mathtt{OptContract}=\mathsf{null}$
while True do
     Find the (approximately) optimal solution to the following program via the algorithm in [Lee et al., 2018], denote the value by $\mathtt{Obj}(k)$:
$\min\sum_{i}\beta_{i},\quad\text{s.t. }f(a(\beta))\geq k$
     if  $(1-\mathtt{Obj}(k))\cdot k\geq\mathtt{OptUtil}$ then
         $\mathtt{OptUtil}=(1-\mathtt{Obj}(k))\cdot k$
         Update $\mathtt{OptContract}$ as the solution to the above program
     end if
     if $\mathtt{Obj}(k)\geq 1$ then
         Break
     end if
     $k\mathrel{\mathop{:}}=k+\varepsilon$
end while
Return $\mathtt{OptContract}$
Define the convex program $\mathtt{MinContract}(k)$ as
$\min\sum_{i}\beta_{i},\quad\text{s.t. }f(a(\beta))\geq k.$
$\mathtt{MinContract}(k)$ returns the minimum possible total share that must be distributed when the production output is required to meet a certain threshold $k$. Denoting the objective as $\mathtt{Obj}(k)$, the following result should be immediate.
Theorem 8.
The optimal utility for the principal is equal to $\sup_{k\in[0,\infty)}(1-\mathtt{Obj}(k))\cdot k$.
Therefore, learning the optimal contract involves solving a family of convex programs. An implementation is given in Algorithm 3, which finds the approximate optimal contract. An especially important aspect to consider is that the convex set in the constraint (i.e., the feasible set $f(a(\beta))\geq k$) is unknown to the learner. This makes popular methods for constrained convex optimization, such as projected gradient descent, unable to be applied. However, the recent work by [Lee et al., 2018] provides algorithms for convex optimization with membership oracles, which exactly suits this setting. [Lee et al., 2018] show that a $\varepsilon$-optimal point can be found using $\widetilde{O}(n^{2})$ queries. Denoting ${k_{\max}}=\max_{\beta\in\Delta(n)}f(a(\beta))$ (here $\Delta(n)$ is the simplex in $\mathbb{R}^{n}$), there are at most $O({k_{\max}}/\varepsilon)$ convex programs that need to be solved, and hence the principal can find the optimal contract using ${k_{\max}}\cdot\widetilde{O}(n^{2}/\varepsilon)$ queries.
8 Concluding Remarks
As a concluding remark, I explain how the one-dimensional effort model helped attain the results in this paper, which suggests this modeling choice may be more appropriate over the discrete choice model in some scenarios.
In the single-agent setting, I attained results for contracting against heterogeneous, homogeneous, and non-myopic agents. For heterogeneous agents, I show how the problem can be directly reduced to Lipschitz bandits under a relatively mild condition. The one-dimensional effort model allowed us to borrow ideas from the first-order approach (e.g. [Rogerson, 1985]), which in turn guaranteed the continuity of the principal’s utility as a function of the contract. By contrast, prior works by [Ho et al., 2014, Zhu et al., 2022] adopted the discrete action space model, under which the principal’s utility will always be discontinuous. For homogeneous (identical) agents, the one-dimensional effort model allowed us to directly use the diminishing returns assumption. This allowed the principal to recover the agent’s production and cost function by only querying linear contracts and gives a polynomial sample complexity scheme, an exponential improvement over prior works (e.g. [Cohen et al., 2022]). For non-myopic agents, under linear contracts and the one-dimensional effort model, the agent’s deviation can be controlled with the idea of delayed feedback ([Haghtalab et al., 2022]), and I give a low strategic-regret mechanism for the principal. Further, I also show a connection between linear contracts in the one-dimensional effort model and posted-price auctions with an unknown demand curve ([Kleinberg and Leighton, 2003]). This connection directly gives a lower bound on the regret for learning the optimal linear contract. I also studied the problem of learning the optimal contract in a team production setting: a condition is identified under which the principal’s learning problem can be reformulated as solving a family of convex programs. I believe this work makes several interdisciplinary advancements at the intersection of online learning and contract design.
Appendix A Missing Proofs for Section 3
Proposition 5.
Under SDFC, for any monotone contract $s$, $u_{\mathsf{A}}(a;s)$ is strictly concave with respect to $a$.
Proof.
Recall the agent’s utility can be expressed as:
$u_{\mathsf{A}}(a;s)$ $=\sum_{j=1}^{m}\Delta s_{j}G_{j}(a)-c(a).$
Taking the second derativie with respect to $a$:
$u^{\prime\prime}_{\mathsf{A}}(a;s)$ $=\sum_{j=1}^{m}\Delta s_{j}G^{\prime\prime}_{j}(a)-c^{\prime\prime}(a)$
$<\sum_{j\in[m]}\Delta s_{j}c^{\prime\prime}(a)-c^{\prime\prime}(a)$
$\leq s_{m}c^{\prime\prime}(a)-c^{\prime\prime}(a)$
$\leq 0.$
This completes the proof. ∎
Lemma 7.
The agent’s response $a(s)$ is Lipschitz continuous with respect to the contract $s$ in the monotone contract domain, specifically, for any two monotone contracts $s^{1},s^{2}$:
|$\absolutevalue{a(s^{1})-a(s^{2})}<\frac{2L}{\lambda-\mu}\norm{s^{1}-s^{2}}_{1}.$
Proof.
Let $a\mathrel{\mathop{:}}=a(s^{1}),b\mathrel{\mathop{:}}=a(s^{2})$ be the best responses. Without loss of generality assume $a<b$.
By the previous Proposition 1, the optimal response must satisfy the first-order conditions.
$\sum_{j\in[m]}s^{1}_{j}f^{\prime}_{j}(a)$ $\leq c^{\prime}(a)$
$\sum_{j\in[m]}s^{2}_{j}f^{\prime}_{j}(b)$ $\geq c^{\prime}(b).$
The above expressed in terms of $G$ and $\Delta s$ becomes:
$\sum_{j\in[m]}\Delta s^{1}_{j}G^{\prime}_{j}(a)$ $\leq c^{\prime}(a)$
$\sum_{j\in[m]}\Delta s^{2}_{j}G^{\prime}_{j}(b)$ $\geq c^{\prime}(b).$
Note that by strong convexity of $c$, the difference between the right-hand side can be lower bounded:
$c^{\prime}(b)-c^{\prime}(a)\geq\lambda(b-a).$
The difference between the left-hand side can be upper bounded:
$\sum_{j\in[m]}\absolutevalue{\Delta s^{1}_{j}G^{\prime}_{j}(a)-\Delta s^{2}_{j}G^{\prime}_{j}(b)}$ | $=\sum_{j\in[m]}\absolutevalue{\Delta s^{1}_{j}G^{\prime}_{j}(a)-\Delta s^{2}_{j}G^{\prime}_{j}(a)-\Delta s^{2}_{j}(G^{\prime}_{j}(b)-G^{\prime}(a))}$ |
[…]
$\leq 2L\norm{s^{1}-s^{2}}_{1}+\mu\absolutevalue{a-b}.$ | .
Therefore, we must have:
[…]
which translates to:
|$\absolutevalue{a-b}<\frac{2L\norm{s^{1}-s^{2}}_{1}}{\lambda-\mu}.$
Hence, when $s$ is restricted to be monotone, the response $a(s)$ is Lipschitz with respect to the contract $s$. ∎
Theorem 9.
As a function of $s$, the utility of the principal $u_{\mathsf{P}}(s)$ is Lipschitz, specifically:
|$\absolutevalue{u_{\mathsf{P}}(s^{1})-u_{\mathsf{P}}(s^{2})}\leq\left(\frac{4L^{2}}{\lambda-\mu}+2\right).\norm{s^{1}-s^{2}}_{1}.$
Proof.
Recall the utility of the principal can be expressed as:
$u_{\mathsf{P}}(s)$ $=\sum_{j}(\pi_{j}-s_{j})f_{j}(a(s))$
$=\sum_{j\in[m]}\pi_{j}f_{j}(a(s))-\sum_{j\in[m]}s_{j}f_{j}(a(s))$
Consider the two terms separately. The first term is the expected outcome, and can be expressed as:
$\sum_{j}\pi_{j}f_{j}(a(s))$ $=\sum_{j\in[m]}(\pi_{j}-\pi_{j-1})G_{j}(a(s))$
which can been seen to be $\frac{2L^{2}}{\lambda-\mu}$-Lipschitz. This follows from the fact that $a(s)$ is $\frac{2L}{\lambda-\mu}$-Lipschitz with respect to $s$, and that $G_{j}(a)$ is $L$ Lipschitz with respect to $a$.
Next, consider the second term $\sum_{j\in[m]}s_{j}f_{j}(a(s))$, which is the expected payment to the agent. Consider two contracts $s^{1},s^{2}$ and let the best response be $a,b$ respectively. Then the difference between the payments can be upper bounded:
|$\absolutevalue{\sum_{j\in[m]}s^{1}_{j}f_{j}(a)-\sum_{j\in[m]}s^{2}_{j}f_{j}(a)}$ | $\leq\sum_{j\in[m]}\absolutevalue{\Delta s^{1}_{j}G_{j}(a)-\Delta s^{2}_{j}G_{j}(b)}$ |
$=\sum_{j\in[m]}\absolutevalue{\Delta s^{1}_{j}G_{j}(a)-\Delta s^{1}_{j}G_{j}(b)-(\Delta s^{2}_{j}-\Delta s^{1}_{j})G_{j}(b)}$ |
[…]
$\leq\left(\frac{2L^{2}}{\lambda-\mu}\right)\norm{s^{1}-s^{2}}_{1}+2\norm{s^{1}-s^{2}}_{1}$
$=\left(\frac{2L^{2}}{\lambda-\mu}+2\right)\norm{s^{1}-s^{2}}_{1}$
In the fourth line, we use the fact that |$\absolutevalue{G_{j}(b)-G_{j}(a)}\leq L\absolutevalue{b-a}$ | and |$\absolutevalue{b-a}\leq(\frac{2L}{\lambda-\mu})\norm{s^{1}-s^{2}}_{1}$ by the previous lemma. This shows the second term is $\left(\frac{2L^{2}}{\lambda-\mu}+2\right)$-Lipschitz.
The utility to the principal is the expected outcome minus payment, hence must be $\left(\frac{4L^{2}}{\lambda-\mu}+2\right)$-Lipschitz. ∎
Appendix B Missing Proofs for Section 4
B.1 The estimated paramters are accurate
We first show the response of the agent is well-behaved under linear contracts.
Lemma 8.
Let $a(\beta)$ denote agent’s best response when the linear contract $\beta$ is posted. Then $a(\beta)$ is increasing and continuous in $\beta$.
Proof.
Recall $r(a)$ is the expected production of the agent. Assume the principal posts a linear contract $\beta$, the agent then takes the action $a=a(\beta)$ that maximizes the following:
$\max_{a}\beta r(a)-c(a).$
Note that this is a concave function in $a$. The optimal solution then satisfies the following.
  1. 1.
    If $\beta<\frac{c^{\prime}(0)}{r^{\prime}(0)}$, the agent’s best response is $a(\beta)=0$,
  2. 2.
    If $\beta>\frac{c^{\prime}(E)}{r^{\prime}(E)}$, the agent’s best response is $a(\beta)=E$,
  3. 3.
    Otherwise, the agent’s best response is the unique actions satisfying:
    $\frac{c^{\prime}(a)}{r^{\prime}(a)}=\beta.$
Note $c^{\prime}$ is increasing, $r^{\prime}$ is decreasing and both $c^{\prime}$ and $r^{\prime}$ are continuous. Hence, the agent’s response is continuous and increasing with respect to $\beta$. The principal is then able to implement every action using linear contracts. ∎
When the contract space is appropriately discretized, the induced actions are also close to each other. In other words, $a(\beta)$ is Lipschitz with respect to $\beta$.
Lemma 9.
Let $\beta^{\prime}>\beta$, then $a(\beta^{\prime})-a(\beta)<\frac{r^{\prime}(0)}{\lambda}(\beta^{\prime}-\beta)$.
Proof.
With out loss of generality let us assume $\frac{c^{\prime}(0)}{r^{\prime}(0)}<\beta<\beta^{\prime}<\frac{c^{\prime}(E)}{r^{\prime}(E)}$. Recall at each linear contract $\beta\in[\frac{c^{\prime}(0)}{r^{\prime}(0)},\frac{c^{\prime}(E)}{r^{\prime}(E)}]$, the induced action $a$ satisfies:
$\frac{c^{\prime}(a)}{r^{\prime}(a)}=\beta.$
Consider $\beta$ as a function of $a$, then
$\beta^{\prime}(a)$ $=c^{\prime}(a)\cdot\frac{-r^{\prime\prime}(a)}{(r^{\prime}(a))^{2}}+c^{\prime\prime}(a)\cdot\frac{1}{r^{\prime}(a)}$
$\geq c^{\prime\prime}(a)\cdot\frac{1}{r^{\prime}(a)}$
$\geq\frac{\lambda}{r^{\prime}(0)}.$
Hence, by the inverse function rule:
$a^{\prime}(\beta)\leq\frac{r^{\prime}(0)}{\lambda}.$
[…]
If only assume $r^{\prime\prime}<\mu$, then $c^{\prime}(a)\cdot-\mu/r^{\prime 2}+\lambda/L$ This is enough to complete the proof. ∎
By querying a specific contract $\beta^{(i)}$ enough times, the principal can learn an accurate estimate of the production technology $f_{j}(a^{(i)})$ (and $G_{j}(a^{(i)})$).
Lemma 10.
Fix the linear contract $\beta^{(i)}\in\mathcal{C}_{D}$. Suppose the contract has been queried for $N=\frac{\ln(2/\delta)}{\varepsilon^{2}}$ times. For any $j\in[m]$, with probability $1-\delta$,
|$\absolutevalue{\hat{G}_{j}(a^{(i)})-G_{j}(a^{(i)})}\leq\varepsilon,$
and
|$\absolutevalue{r(a^{(i)})-\hat{r}(a^{(i)})}\leq\varepsilon.$
Proof.
Follows directly from the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality. ∎
There is another question we need to solve: how can the principal infer the cost of the agent’s action under some contract $\beta$? This can be solved using an adaptation of Myerson’s lemma.
Lemma 11.
Let $a=a(\beta)$ be the action induced by $\beta$, $a^{\prime}=a(\beta^{\prime})$ be the action induced by $\beta^{\prime}$. Then the following holds:
$\beta^{\prime}[r(a)-r(a^{\prime})]\leq c(a)-c(a^{\prime})\leq\beta[r(a)-r(a^{\prime})]$
Proof.
By incentive compatibility:
$\beta r(a)-c(a)$ $\geq\beta r(a^{\prime})-c(a^{\prime})$
$\beta^{\prime}r(a^{\prime})-c(a^{\prime})$ $\geq\beta^{\prime}r(a)-c(a)$
The lemma then follows by rearranging terms. ∎
The below lemma characterizes the confidence interval for the cost of the action $a(\beta^{(i)})$.
Lemma 12.
Let
${c}^{\mathtt{LCB}}(a^{(i)})=\sum_{j=1}^{i}\beta^{(j-1)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))-3\varepsilon,$
${c}^{\mathtt{UCB}}(a^{(i)})=\sum_{j=1}^{i}\beta^{(j)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))+3\varepsilon$
be as defined in the algorithm. Then for any $i$
$c(a^{(i)})\in\left[{c}^{\mathtt{LCB}}(a^{(i)}),{c}^{\mathtt{UCB}}(a^{(i)})\right].$
Further
${c}^{\mathtt{UCB}}(a^{(i)})-{c}^{\mathtt{LCB}}(a^{(i)})\leq 9\varepsilon.$
Proof.
For any $i\geq 1$, the following holds.
$c(a^{(i)})$ $=\sum_{j=1}^{i}c(a^{(j)})-c(a^{(j-1)})$
$\leq\sum_{j=1}^{i}\beta^{(j)}(r(a^{(j)})-r(a^{(j-1)}))$
$=\beta^{(i)}r(a^{(i)})+\sum_{j=2}^{i}(\beta^{(j-1)}-\beta^{(j)})r(a^{(j-1)})-\beta_{1}r(a^{(0)})$
$\leq\beta^{(j)}\hat{r}(a^{(j)})+\sum_{j=2}^{i}(\beta^{(j-1)}-\beta^{(j)})\hat{r}(a^{(j-1)})-\beta_{1}\hat{r}(a^{(0)})+3\varepsilon$
$\leq\sum_{j=1}^{i}\beta^{(j)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))+3\varepsilon.$
Similarly we can obtain a lower bound for $c(a^{(i)})$.
$c(a^{(i)})$ $=\sum_{j=1}^{i}c(a^{(j)})-c(a^{(j-1)})$
$\geq\sum_{j=1}^{i}\beta^{(j-1)}(r(a^{(j)})-r(a^{(j-1)}))$
$=\beta^{(i-1)}r(a^{(i)})+\sum_{j=2}^{i}(\beta^{(j-2)}-\beta^{(j-1)})r(a^{(j-1)})-\beta^{(0)}r(a^{(0)})$
$\geq\beta^{(i-1)}\hat{r}(a^{(i)})+\sum_{j=2}^{i}(\beta^{(j-2)}-\beta^{(j-1)})\hat{r}(a^{(j-1)})\beta^{(0)}-\beta^{(0)}\hat{r}(a^{(0)})-3\varepsilon$
$\geq\sum_{j=1}^{i}\beta^{(j-1)}(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))-3\varepsilon.$
Further,
${c}^{\mathtt{UCB}}(a^{(i)})-{c}^{\mathtt{LCB}}(a^{(i)})$ $\leq\sum_{j=1}^{i}(\beta^{(j)}-\beta^{(j-1)})(\hat{r}(a^{(j)})-\hat{r}(a^{(j-1)}))+6\varepsilon$
$\leq\sum_{j=1}^{i}(\beta^{(j)}-\beta^{(j-1)})3\varepsilon+6\varepsilon$
$\leq 3\varepsilon+6\varepsilon$
$\leq 9\varepsilon,$
where in the second line we use the fact that:
|$\absolutevalue{\hat{r}(\beta^{(j)})-\hat{r}(\beta^{(j-1)})}$ | $\leq 2\varepsilon+r(\beta^{(j)})-r(\beta^{(j-1)})$
$\leq 2\varepsilon+r^{\prime}(0)\cdot(a(\beta^{(j)})-a(\beta^{(j-1)}))$
$\leq 2\varepsilon+r^{\prime}(0)\cdot\frac{r^{\prime}(0)}{\lambda}\cdot\varepsilon_{c}$
$\leq 3\varepsilon.$
This finishes the proof. ∎
B.2 The family of linear programs gives an approximately optimal contract
Recall for each $i\in\mathinner{\left\{0,1,\dots,\left\lfloor 1/\varepsilon_{c}\right\rfloor\right\}}$, the principal solves the following linear program, denote the program by $\mathtt{LP}(a^{(i)})$ and its value by $\mathtt{APX}(a^{(i)})$.
$\max_{s}$ $\quad\hat{r}(a^{(i)})-\sum_{j\in[m]}s_{j}\hat{f}_{j}(a^{(i)})$
$u^{\mathtt{UCB}}_{\mathsf{A}}(a^{(i)};s)$ $\geq u^{\mathtt{LCB}}_{\mathsf{A}}(a^{(k)};s)-O(\varepsilon)\quad\forall a^{(k)}\in\mathcal{A}_{D}$
$s_{j}$ $\geq s_{j-1}$
$s_{j}$ $\geq 0$
Lemma 13.
Fix any action $a$ and any monotone bounded contract $s$. Let $\widebar{a}$ be the action in $\mathcal{A}_{D}$ that is closest to $a$. Then,
$u_{\mathsf{A}}(a;s)-u_{\mathsf{A}}(\widebar{a};s)\leq\varepsilon.$
Proof.
We have
|$\absolutevalue{u_{\mathsf{A}}(a;s)-u_{\mathsf{A}}(\widebar{a};s)}$ | |$\leq\absolutevalue{\sum_{j\in[m]}s_{j}[f_{j}(a)-f_{j}(\widebar{a})]}+\absolutevalue{c(a)-c(\widebar{a})}$ |
|$\leq\absolutevalue{\sum_{j\in[m]}\Delta s_{j}[G_{j}(a)-G_{j}(\widebar{a})]}+\absolutevalue{c(a)-c(\widebar{a})}$ |
$\leq\sum_{j\in[m]}\Delta s_{j}\cdot L\cdot\absolutevalue{a-\widebar{a}}+c^{\prime}(E)\cdot\absolutevalue{a-\widebar{a}}$ |
$\leq(L+c^{\prime}(E))\cdot\absolutevalue{a-\widebar{a}}$ |
$\leq(L+c^{\prime}(E))\cdot\frac{r^{\prime}(0)}{\lambda}\varepsilon_{c}$
$\leq\frac{2L^{2}}{\lambda}\varepsilon_{c}$
$\leq\varepsilon.$
Here in the last line, we used the fact that:
$\varepsilon_{c}\leq\frac{1}{L+c^{\prime}(E)}\cdot\frac{\lambda}{r^{\prime}(0)}.$
[…]
Need condition here again. ∎
Lemma 14.
$0\leq u^{\mathtt{UCB}}(a^{(i)};s)-u_{\mathsf{A}}(a^{(i)};s)\leq 10\varepsilon$, $0\leq u_{\mathsf{A}}(a^{(i)};s)-u^{\mathtt{LCB}}_{\mathsf{A}}(a^{(i)};s)\leq 10\varepsilon$.
Proof.
$u^{\mathtt{UCB}}(a^{(i)};s)-u_{\mathsf{A}}(a^{(i)};s)$ $=\sum_{j\in[m]}\Delta s_{j}[G^{\mathtt{UCB}}_{j}(a^{(i)})-G_{j}(a^{(i)})]+c(a^{(i)})-c^{\mathtt{LCB}}(a^{(i)})$
$\leq 10\varepsilon.$
And similarly for others. ∎
We now show $a^{(i)}$ will be $\delta$-IC under any solution to the program $\mathtt{LP}(a^{(i)})$.
Lemma 15.
Consider the program $\mathtt{LP}(a^{(i)})$ and let $s$ be a solution. Then $a^{(i)}$ is $\delta$-IC under $s$ with $\delta=22\varepsilon$.
Proof.
Fix any action $a$, then by the previous Lemma 13, there exists some $\widebar{a}\in\mathcal{A}_{D}$ such that
$u_{\mathsf{A}}(a;s)\leq u_{\mathsf{A}}(\widebar{a};s)+\varepsilon.$
Further:
$u_{\mathsf{A}}(a^{(i)};s)$ $\geq u^{\mathtt{UCB}}_{\mathsf{A}}(a^{(i)};s)-10\varepsilon$
$\geq u^{\mathtt{LCB}}_{\mathsf{A}}(\widebar{a};s)-11\varepsilon$
$\geq u_{\mathsf{A}}(\widebar{a};s)-21\varepsilon$
$\geq u_{\mathsf{A}}(a;s)-22\varepsilon.$
Here, the second line follows from the fact that $a^{(k)}$ satisfies the (approximate) incentive compatibility constraint in $\mathtt{LP}(a^{(i)})$, and the fourth line follows from the fact that $a$ and $\widebar{a}$ are close and applying lemma 13. This shows $a^{(i)}$ is indeed $\delta$-IC with $\delta=22\varepsilon$. ∎
Next, let us relate the utility of the optimal contract to the contract computed by the estimated parameters. Let us consider the true optimal contract $s^{*}$, and denote the induced action under $s^{*}$ by $a^{*}$. Let us denote $\widebar{a^{*}}$ to be the action in $\mathcal{A}_{D}$ that is closest to $a^{*}$.
Lemma 16.
$s^{*}$ is a feasible solution to $\mathtt{LP}(\widebar{a^{*}})$.
Proof.
That $s^{*}$ is monotone is clear. We only need to show $s^{*}$ satisfies the approximate-IC constraint. Fix any $a^{(k)}\in\mathcal{A}_{D}$. Note:
$u_{\mathsf{A}}^{\mathtt{UCB}}(\widebar{a^{*}};s^{*})$ $\geq u_{\mathsf{A}}(\widebar{a^{*}};s^{*})$
$\geq u_{\mathsf{A}}(a^{*};s^{*})-\varepsilon$
$\geq u_{\mathsf{A}}(a^{(k)};s^{*})-\varepsilon$
$\geq u_{\mathsf{A}}^{\mathtt{LCB}}(a^{(k)};s^{*})-\varepsilon.$
Here, the second line is because $a^{*}$ and $\widebar{a^{*}}$ are close and applying lemma 13, the third line is because $a^{*}$ is incentive compatibly under $s^{*}$. ∎
Lemma 17.
Fix any monotone contract $s$. $\mathtt{APX}(a^{(i)},s)\leq u_{\mathsf{P}}(a^{(i)};s)+2\varepsilon$.
Proof.
$\mathtt{APX}(a^{(i)};s)-u_{\mathsf{P}}(a^{(i)};s)=\hat{r}(a^{(i)})-r(a^{(i)})+\sum_{j\in[m]}s_{j}[f_{j}(a^{*})-\hat{f}_{j}(a^{*})]$
$\sum_{j\in[m]}s_{j}[f_{j}(a^{*})-\hat{f}_{j}(a^{*})]$ $\leq\sum_{j\in[m]}\Delta s_{j}[G_{j}(a^{*})-\hat{G}_{j}(a^{*})]$
$\leq\sum_{j\in[m]}\Delta s_{j}\cdot\varepsilon$
$\leq\varepsilon.$
Further $\hat{r}(a^{(i)})-r(a^{(i)})<\varepsilon$. ∎
Let $a^{*}_{D}$ be the action with the highest objective among the family of $\mathtt{LP}(a^{(i)})$, and let $s^{*}_{D}$ be the solution to $\mathtt{LP}(a^{*}_{D})$.
Proposition 6.
Under the contract $s_{D}^{*}$, the action $a_{D}^{*}$ is $\delta$-IC with $\delta=22\varepsilon$. Moreover, $\mathtt{OPT}\leq u_{\mathsf{P}}(a^{*}_{D};s^{*}_{D})+6\varepsilon$.
Proof.
For the first part, that $a_{D}^{*}$ is $\delta$-IC follows from lemma 15. For the second part,
$\mathtt{OPT}$ $=r(a^{*})-\sum_{j\in[m]}s^{*}_{j}f_{j}(a^{*})$
$\leq r(\widebar{a^{*}})-\sum_{j\in[m]}s^{*}_{j}f_{j}(\widebar{a^{*}})+2\varepsilon$
$\leq\mathtt{APX}(\widebar{a^{*}};s^{*})+4\varepsilon$
$\leq\mathtt{APX}(a^{*}_{D};s^{*}_{D})+4\varepsilon$
$\leq u_{\mathsf{P}}(a^{*}_{D};s^{*}_{D})+6\varepsilon.$
Here, the second line follows from the fact that $a^{*}$ and $\widebar{a^{*}}$ are close, the third line follows from the previous lemma 17, the fourth line follows from the fact that $s_{D}^{*}$ is optimal among the family of LPs using estimated parameters, and the final line is from lemma 17 again. ∎
Now, we apply the very interesting observation in [Dütting et al., 2021], which relates the utility of $\delta$-IC contracts to that of exact IC contracts. For completeness, we include the proof (phrased in a slightly different way from their original proof).
Lemma 18 ([Dütting et al., 2021]).
Let $(s,a)$ be a $\delta$-IC contract. Then $s^{\prime}\mathrel{\mathop{:}}=(1-\sqrt{\delta})s+\sqrt{\delta}\pi$ achieves utility at least $(1-\sqrt{\delta})u_{\mathsf{P}}(a;s)-(\sqrt{\delta}-\delta)$.
Proof.
Denote $p(a;s)$ to be the expected payment under $s$ if the agent were to take action $a$. It is easy to calculate
$u_{\mathsf{P}}(a;s^{\prime})=(1-\sqrt{\delta})u_{\mathsf{P}}(a;s)$
Now consider the contract $s^{\prime}$, and suppose this contract implements $a^{\prime}$, then IC is satisfied:
$p(a^{\prime};s^{\prime})-c(a^{\prime})\geq p(a;s^{\prime})-c(a),$
a direct calculation gives:
$(1-\sqrt{\delta})p(a^{\prime};s)+\sqrt{\delta}r(a^{\prime})-c(a^{\prime})\geq(1-\sqrt{\delta})p(a;s)+\sqrt{\delta}r(a)-c(a).$
Simplifying further:
$p(a^{\prime};s)-c(a^{\prime})+\sqrt{\delta}[r(a^{\prime})-p(a^{\prime};s)]\geq p(a;s)-c(a)+\sqrt{\delta}[r(a)-p(a;s)]$
Since $a$ is $\delta$-IC under $s$, then
$p(a;s)-c(a)\geq p(a^{\prime};s)-c(a^{\prime})-\delta.$
The above two inequalities directly gives:
$\sqrt{\delta}[r(a^{\prime})-p(a^{\prime};s)]\geq\sqrt{\delta}[r(a)-p(a;s)]-\delta$
The principal’s utility under the exact IC contract $s^{\prime}$ can be lower bounded:
$u_{\mathsf{P}}(a^{\prime};s^{\prime})$ $=(1-\sqrt{\delta})(r(a^{\prime})-p(a^{\prime};s))$
$\geq(1-\sqrt{\delta})(r(a)-p(a;s)-\sqrt{\delta})$
$\geq(1-\sqrt{\delta})(r(a)-p(a;s))-(\sqrt{\delta}-\delta).$
This completes the proof. ∎
Theorem 10.
The principal can find a $6\varepsilon+10\sqrt{\varepsilon}$ approximately optimal contract using $\widetilde{O}(1/\varepsilon^{3})$ samples.
Proof.
Recall the contract $(s_{D}^{*},a_{D}^{*})$ is $22\varepsilon$-IC and the utility is at least a $6\varepsilon$ approximation. Then, using the above result, $s_{D}^{*}$ can be converted into an exact IC contract that achieves utility at least:
$(1-\sqrt{22\varepsilon})(u_{\mathsf{P}}(a^{*};s^{*})-6\varepsilon)-(\sqrt{22\varepsilon}-22\varepsilon)$ $\geq u_{\mathsf{P}}(a^{*};s^{*})-\sqrt{22\varepsilon}u_{\mathsf{P}}(a^{*};s^{*})-6\varepsilon-\sqrt{22\varepsilon}$
$\geq u_{\mathsf{P}}(a^{*};s^{*})-6\varepsilon-10\sqrt{\varepsilon}.$
[…]
B.3 Can assume $\beta=1$ induce maximum action
Suppose $\beta=1$ induces the action $a<E$. Let $\beta=1$ induce the action $a$. We show any action $b>a$ will be too costly to induce. Let us suppose the contract $s$, induces $b$. Then
$\sum_{j}s_{j}f_{j}(b)-c(b)>\sum_{j}s_{j}f_{j}(a)-c(a).$
Also, since the linear contract $\beta=1$ induces $a$, then
$\sum_{j}r(a)-c(a)>\sum_{j}r(b)-c(b).$
Then:
$\sum_{j}(\pi_{j}-s_{j})f_{j}(a)>\sum_{j}(\pi_{j}-s_{j})f_{j}(b)$
So the question is, does there exist a positive sequence such that:
$\sum_{j}y_{j}(f_{j}(a)-f_{j}(b))>0$
B.3.1 Idea 2
Let us suppose that contract $s$ induce $b$. Then:
$\sum_{j}\Delta s_{j}G_{j}(b)-c(b)>\sum_{j}\Delta s_{j}G_{j}(a)-c(a)$
Also:
$r(a)-c(a)>r(b)-c(b)$
The utility of inducing $b$ is:
$r(b)-\sum_{j}\Delta s_{j}G_{j}(b)<c(b)-c(a)-\sum_{j}\Delta s_{j}G_{j}(a)$
B.3.2 Idea 3
By first order condition.
$\sum_{j}s_{j}G^{\prime}_{j}(b)$ $=c^{\prime}(b)$
$\sum_{j}\pi_{j}G^{\prime}_{j}(b)$ $<c^{\prime}(b)$
Second line comes from $r^{\prime}(b)<c^{\prime}(b)$.